OpenStudy (anonymous):

x^3 TIMES x^2 Is it just x^5?

zepdrix (zepdrix):

\(\large\rm x^a\cdot x^b=x^{a+b}\) yes, good :)

OpenStudy (anonymous):

Yes. Every time you multiply exponents you multiply the base number and just add the exponents together. Another example is 4^3*5^4=20^7

jimthompson5910 (jim_thompson5910):

`4^3*5^4=20^7` is not true in order to add the exponents, the bases must be the same. Zepdrix has the right idea

OpenStudy (anonymous):

Oh yeah you're right, wow I'm tired...Been doing math all day, lol sorry ;) The sad part is I know this very well too